OpenStudy (anonymous):

Are there values of c and m that make { cx^2 if x <1 h(x) = { 4 if x =1 {-x^2 + mx if x >1 continuous at x +1? Find c and m if they exist, or explain why they do not exists

OpenStudy (anonymous):

this is not as hard as it looks put \(x=1\) in the first two expressions and solve i.e. solve \[c\times 1^2=4\] which pretty clearly makes \(c=4\)

OpenStudy (anonymous):

typo for the second one i meant solve \[-1^2+m\times 1=4\]

OpenStudy (anonymous):

But i thought they were asking at x+1?

OpenStudy (anonymous):

@@satellite73 Is it different at x +1 or at x? On my paper here it is "at x +1 " Does the question have error or somethin?

OpenStudy (aum):

Looks like a typo. The "+" symbol and "=" symbol are on the same key and they must have pressed the shift key by mistake. "continuous at x = 1" makes sense here but "continuous at x +1" does not.